An open container holds ice of mass 0.500 kg at a temperature of -10.8 ° C. The mass of the container can be ignored. Heat is supplied to the container at the constant rate of 800 J/minute The specific heat of ice to is 2100 J/kg K and the heat of fusion for ice is 334 x 103 J/kg.
How much time tmelts passes before the ice starts to melt?
Answer #1

answer) we know the formula

Q=mcdT=0.500kg*2100J/kg.K*(0-(-10.8)=11340J

so t melt=Q/800J/min=11340J/800J/min=14.175 minutes

so the answer is 14 minutes or 14.2 minutes or 14.18 minutes or 14.175 minutes